OpenStudy (anonymous):

please help! :) Write in standard form: y²+4y-8x+4=0

OpenStudy (theviper):

wait :)

OpenStudy (theviper):

\[\large{ y²+4y-8x+4=0}\]\[\large{=>y^2+4y-8x=-4}\]

OpenStudy (theviper):

got it??

OpenStudy (anonymous):

thanks! is that standard form? somebody else told me this: 8x=(y+2)² which is correct? @TheViper

OpenStudy (theviper):

mine :) well I wanna check because I just understand to write an equation in standard form :) @ash2326 plz help :)

OpenStudy (theviper):

@Hero :)

OpenStudy (theviper):

Can u tell me I did write or wrong ??

hero (hero):

\[(y+2)^2 - 8x = 0\]

OpenStudy (theviper):

Yes I did not solve that further :) So @erin512 @Hero is correct :) Sure ;)

OpenStudy (theviper):

Got it ?

OpenStudy (theviper):

@Hero hates math in picture otherwise he loves math :D

OpenStudy (anonymous):

@Hero @TheViper thank you so much :)

OpenStudy (theviper):

yw :)